pi/2 ∫ 0 dx/1 + tan x

pi/20 dx/1 + tan x

Grade:12

2 Answers

abhishek singh
9 years ago
answer will be pi/4 ..see   we have to use the property of definite integral f(a-x) = f(x).
   also change tanx to sinx / cosx 
 
 now we have  pi/20 dx/1 + tan x =  pi/20 dx/ (1 + (sinx/cosx) )  = pi/20 (cosx)dx /(sinx + cosx)  
 now apply    f(a-x) = f(x).   therfore we get ½  pi/20 (cosx + sinx )dx /(sinx + cosx)    =    1/2*  [1]pi/2       = pi/4
                                                                                                                                                  0
abhishek singh
9 years ago
sorrythelast line was wrongly  typed  ..  it was lower limit = 0  and upper limit = pi/2  of [1]* 1/2